Explain how light is emitted when an electron is excited
1
Expert's answer
2020-10-07T07:32:50-0400

Solution.

When an atom passes from the excited state to the ground state, the electron passes from the (n+1) level to the n-th level. And during this process, energy is released, namely in the form of light.

"En = -\frac{me^4}{8h^2e0^2n^2}"

"\Delta E = -\frac{me^4}{8h^2e0^2}(\frac{1}{n2^2} -\frac{1}{n1^2})"

Answer:

When an atom passes from the excited state to the ground state, the electron passes from the (n+1) level to the n-th level. And during this process, energy is released, namely in the form of light.